OpenStudy (anonymous):

Identify the first step using the elimination method to solve the following system of equations: 10x - 7y = -18 15x + 3y = 27 A. 7(10x - 7y = -18) 3(15x + 3y = 27) B. 15(10x - 7y = -18) 10(15x + 3y = 27) C. -3(10x - 7y = -18) 2(15x + 3y = 27) D. 27(10x - 7y = -18) 18(15x + 3y = 2

OpenStudy (amorfide):

first step for elimination, is to get the coefficient of x, or coefficient of y to be the same value

OpenStudy (amorfide):

look through your given answers, multiply out the brackets and see which one gives you a coefficient of x to be the same in both equations, or the same coefficient of y to be the same

OpenStudy (amorfide):

if I had 2x+y=3 x+3y=4 I would multiply my second equation by 2 to get the same x coefficient which would give me 2x+6y=8